# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

### v2

@@ Abstract.
In Australia, post-contact language varieties, including creoles and local varieties of international languages, emerged as a result of forced contact between Indigenous communities and English speakers. These contact varieties are widely used, yet are poorly supported by language technologies. This gap presents barriers to participation in civil and economic society for Indigenous communities using these varieties, and reproduces minoritisation of contemporary Indigenous sociolinguistic identities. This paper concerns three questions regarding this context. First, can speech technologies support speakers of Australian Aboriginal English, a local indigenised variety of English? Second, what risks are inherent in such a project? Third, what technology development practices are appropriate for this context, and how can researchers integrate meaningful community participation in order to mitigate risks? We argue that opportunities do exist—as well as risks—and demonstrate this through a case study exploring design practices in a real-world project aiming to improve speech technologies for Australian Aboriginal English. We discuss how we integrated culturally appropriate and participatory processes throughout the project. We call for increased support for languages used by Indigenous communities, including contact varieties, which provide practical economic and socio-cultural benefits, provided that participatory and culturally safe practices are enacted.
@@ 1. Introduction
Most work on speech and language technologies, a.k.a. Natural Language Processing (NLP), focuses on languages which are spoken by majority populations and heavily standardised within one or more countries (Bender, 2009; Sawicki et al., 2023; Joshi et al., 2020). The field of NLP is thus reproducing the minoritisation of language varieties, and in the process reinforcing barriers of exclusion for diverse language communities. These include most local languages in highly linguistically diverse countries such as Indonesia (Aji et al., 2022), as well as various creoles which combine elements of settler-colonial languages elsewhere in the world (Lent et al., 2021). Additionally affected are linguistic varieties of national languages spoken by minoritised groups (Kantharuban et al., 2023), such as Māori English in New Zealand, Multicultural London English in the UK, and African-American English in the US. Speakers of minoritised varieties must typically switch to a majority variety when using language technologies, and when they do they can experience negative psychological impacts, such as feeling that technology wasn’t designed for them (Cunningham et al., 2024; Mengesha et al., 2021; Harrington et al., 2022; Wenzel et al., 2023).
@@
Machine learning-driven NLP is not typically oriented towards participatory processes, driven instead by the valorisation of efficiency (Birhane et al., 2022b) and fundamentally constrained by tensions between participation and scale (Young et al., 2024). NLP evaluations emphasise quantitative and decontextualised metrics (Hutchinson et al., 2022), while language dataset projects often reduce participation to crowdsourced platform work. These projects fail to create relationships which emphasise reciprocity and social justice (Sloane et al., 2022; Cooper and Zafiroglu, 2024). Participatory AI projects, on the other hand, are challenged by a lack of shared understanding of good practices (Groves et al., 2023), and may risk “participation-washing” (Sloane et al., 2022). More generally, projects which aspire to participation often fail to afford decision-making power to participating communities (Arnstein, 1969), and risk co-optation when delivered through corporate structures (Asaro, 2000; Birhane et al., 2022a).
@@
In Australia, where this research takes place, of the 400+ ancestral Indigenous languages that existed before European colonisation (Bowern, 2023, p. 56), most are either extinct or critically endangered (Australian Government et al., 2020). This is due to linguicidal practices of colonialism, which at various times included governments forcibly separating Aboriginal children from their families and coercing Aboriginal pupils to speak English while forbidding them from speaking their languages (Rademaker, 2018). As a result, most Indigenous people in Australia speak Aboriginal English, making it a strong encoder of ethno-cultural identity (Rodríguez Louro and Collard, 2024). This indigenised variety of English integrates phonological, lexico-grammatical and discourse-pragmatic features from English and Aboriginal ancestral languages (see §2.1).
@@
To address these questions, this paper contributes the first detailed consideration of language technologies to support Aboriginal English speakers. To our knowledge, this is the first paper to focus on technologies for this language community. We do so through a case study of an ongoing project intended to improve Automatic Speech Recognition (ASR) for Indigenous communities in Australia. In doing so, this paper follows previous NLP literature in considering a specific Indigenous language community and its context (e.g., Zhang et al., 2022; Teodorescu et al., 2022), with a focus on design considerations in order to avoid assuming that mainstream methods and goals are locally applicable (Noe and Kirshenbaum, 2024; Pine et al., 2024). As such, this paper contributes to our understanding of ethical and responsible development of technologies for minoritised communities, as well as to our understanding of how to design and build culturally sensitive technologies while incorporating participatory methods.
The paper begins with an introduction to Australian Aboriginal English, including a description of relevant linguistic and sociolinguistic factors (§2). We also present a brief survey of related work on Indigenous language technologies. Next, we discuss the opportunities and risks inherent in deploying ASR systems for this community, as well as opportunities and risks in projects building such technologies (§3). Building on this analysis, we explore strategies to mitigate risk, and facilitate opportunities, through a detailed discussion of the design considerations of a real-world project to design culturally appropriate ASR while integrating participatory methods (§4). In doing so, we provide a detailed discussion of some of the considerations that are critical for a project of this sort, including system requirements, data elicitation and transcription, system evaluation, and processes and frameworks for participation, and employment. Finally, we discuss further technological opportunities for supporting speakers of Aboriginal English (§5).
@@ 2. Background and Related Work
When considering how technologies can support speakers of a specific Indigenous language variety, it is critical to have some basic knowledge of the speech community, the language variety, and their contexts (Zhang et al., 2022). In this section we introduce Aboriginal English and its context, followed by a brief summary of prior work relevant to language technologies for Indigenous communities. This includes consideration of both ancestral and contact language varieties, and also of participatory methods in Indigenous language projects.
@@ 2.1. Australian Aboriginal English
Aboriginal English is a contact variety of English used by Indigenous people in Australia (Rodríguez Louro and Collard, 2021). It is particularly prevalent across southern Australia where ancestral languages are less widely spoken. In these contexts, Aboriginal English is used to strengthen Indigenous ways of being and communicating (Eades, 2012a). Aboriginal English has its origins in the product of forced contact between Indigenous people and the English-speaking settler-colonialists. The English spoken by the British was a powerful instrument used to strip Indigenous communities of their ancestral rights (Malcolm, 2000, p. 125). Aboriginal English’s precursor was a mixed jargon viewed as ‘broken English’ by the British. By the time this jargon had transitioned to a novel pidgin (known as ‘New South Wales pidgin’ (Troy, 1994)), it began to travel outside present-day Sydney and across New South Wales. With the advent of the pastoral industry, the pidgin also spread north into Queensland and the Northern Territory, giving rise to Queensland Pidgin English and the Northern Australian creoles, including Roper River Kriol (Meakins, 2014, p. 371).
@@
The Aboriginal English has linguistic characteristics, including pronunciation, lexis, grammar, and discourse-pragmatics, that distinguish it from mainstream English (Eades, 2012a). Pronunciation features include: frequent /h/-dropping and /h/-addition; metathesis of ‘k’ and ‘s’ in the word ‘ask’, which is pronounced ‘aks’; and default reduction of the ‘ing’ verbal suffix to ‘in’. Grammatical features include double negatives as in ‘never done nothing’; and double subjects as in ‘My niece, she did it.’ Lexical features include lexical transfer such as blackfella, meaning ‘black person/fellow’, and borrowings from ancestral languages, such as dardy ‘good/good looking’ from Nyungar.
@@
Experts remain ambivalent as to whether Aboriginal English represents a single ethnolect, or whether different varieties of Aboriginal English, or ‘Aboriginal Englishes’, exist across Australia (Dickson, 2019). Aboriginal English is frequently placed on a continuum that ranges from ‘lighter’ acrolectal varieties (closer to standardised Australian English) to ‘heavier’ basilectal varieties which are closer to Kriol, spoken across northern Australia (Schultze-Berndt et al., 2013). The basilectal varieties are mostly unintelligible to standardised English speakers, but the acrolectal varieties are superficially similar to Australian English.
@@
Because Aboriginal English is English based, it is sometimes socially stigmatised, erroneously classified by laypeople as an incorrect version of English. As a result, L1 users of Aboriginal English are often discriminated against in schooling and legal settings (Harkins, 1994; Eades, 2012b), perpetuating invisibility of the variety and leading to the false perception that its speakers are uneducated and rude (Eades, 2013). Stigmatisation of Aboriginal English is evident in the comment below, which was offered in response to an online article on the role of Aboriginal English in Indigenous health communication (Rodríguez Louro and Collard, 2021).
@@
The commenter’s deliberate use of ‘Pidgin English’ reflects a deficit perspective that undermines the value of Aboriginal English. Rather than embracing Aboriginal English as a linguistic variety many Indigenous people use daily in Australia, this perspective portrays racialised peoples’ language varieties as ‘lacking’ vis-à-vis mainstream English (Hudley et al., 2020, p. e214).
These negative attitudes are exacerbated because Aboriginal English is primarily oral, with no nationally adopted guidelines as to how it should be written. Additionally, the orthographic representation of Aboriginal English varies across locales, reflecting the local linguistic biographies and practices of distinct Indigenous groups. For example, the name of the Traditional Owners of the region comprising the southwest of Western Australia, as well as their language, is variably spelled ‘Nyungar’, ‘Noongar’, or ‘Nyoongar’ depending on community members’ preferences. [^1] Collard is Nyungar so we prefer this spelling. Rather than a weakness, the lack of standardising orthographies for Aboriginal English reflects Indigenous people’s prerogative to differentiate—rather than homogenise—how Aboriginal English is used across Indigenous Country.
Indigenous people tend to find the use of direct questions awkward or offensive (Rodríguez Louro and Collard, 2024), [^2] With some exceptions, including teaching contexts where questions instructors asking questions is heavily routinised (Moses et al., 2008). and these pragmatic considerations continue in Aboriginal English. Direct questions are also likely to elicit misleading information; Indigenous people have been known to answer ‘yes’ or agree with non-Indigenous people even if the question is not fully understood—a phenomenon known as ‘gratuitous concurrence’ (Liberman, 1985, p. 137). To avoid these limitations, researchers have used ‘yarning’, an Indigenous cultural form of storytelling and conversation, to lead unstructured data collection with Indigenous communities who speak Aboriginal English (rodriguez2021working).
@@ 2.3. Technologies for contact language varieties
The NLP literature has to-date had marginal concern with supporting Indigenous contact language communities, such as Hawaiian Pidgin, Australian Kriol, Tok Pisin, Chavacano, etc. For example, Nigerian Pidgin has over 100 million speakers (Lin et al., 2024), yet is poorly supported by technology. This seems to be due to a combination of reasons, including the stigmatisation of contact languages (Siegel, 1999; Lent et al., 2024) and a belief that ancestral languages are more important in terms of global heritage. However Indigenous peoples are not museum artifacts or relics of an ‘ancient’ era (cf.  Gaby and Woods, 2020, e274); their cultures and societies are dynamic and evolving.
@@ 2.4. Culture, and participatory methods, in language research
Recently, the NLP community has demonstrated increasing interest in the topic of culture, with, e.g., two recent workshops on Cross-cultural Considerations in NLP (Dev et al., 2023; Prabhakaran et al., 2024). This interest has been spurred by the capabilities of Large Language Models (LLMs) with regards to linguistic form, while also demonstrating lack of cultural nuance (Hershcovich et al., 2022; Naous et al., 2024; Li et al., 2024). Threads of research have focused on measuring cultural biases (Naous et al., 2024; Tao et al., 2024) and expression of cultural values by LLMs (Arora et al., 2023; Durmus et al., 2023; Ramezani and Xu, 2023).
@@
Within sociolinguistics, the ‘yarning’ method has demonstrated the effectiveness of culturally appropriate, participatory research methods when working with Aboriginal communities on speech data collection projects (Anonymous, YYYY). ‘Yarning’ favours group-based unstructured storytelling over individual interviews, which aligns with the Indigenous cultural protocol of centring the group over the individual (Malcolm, 2018, p. 17). Engaging in groups makes people feel less self-conscious in speech data collection activities than in one-on-one interactions with a researcher. Previous research led by Collard, who is Indigenous, demonstrates how using such culturally appropriate, participatory methods throughout the research process—from choosing data collection protocols and fieldwork location to participant choice—enables recording of the authentic voices of people infrequently represented in academic research.
@@ 2.5. Projects with Indigenous communities
Indigenous varieties are typically differentiated functionally within a multilingual community, and technologies may be more useful for some contact varieties than others (Doğruöz and Sitaram, 2022; Bird, 2022; Lent et al., 2022). Technologists should understand the technology needs of Indigenous communities, which may be distinct from those of non-Indigenous users (Liu et al., 2022; Cooper et al., 2024).

### v3

@@ Abstract.
In Australia, post-contact language varieties, including creoles and local varieties of international languages, emerged as a result of forced contact between Indigenous communities and English speakers. These contact varieties are widely used, yet are poorly supported by language technologies. This gap presents barriers to participation in civil and economic society for Indigenous communities using these varieties, and reproduces minoritisation of contemporary Indigenous sociolinguistic identities. This paper concerns three questions regarding this context. First, can speech technologies support speakers of Australian Aboriginal English, a local indigenised variety of English? Second, what risks are inherent in such a project? Third, what technology development practices are appropriate for this context, and how can researchers integrate meaningful community participation in order to mitigate risks? We argue that opportunities do exist—as well as risks—and demonstrate this through a case study exploring design practices in a project aiming to improve speech technologies for Australian Aboriginal English. We discuss how we integrated culturally appropriate and participatory processes throughout the project. We call for increased support for languages used by Indigenous communities, including contact varieties, which provide practical economic and socio-cultural benefits, provided that participatory and culturally safe practices are enacted.
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732010isbn: 979-8-4007-1482-5/2025/06
@@ 1. Introduction
Most work on speech and language technologies, a.k.a. Natural Language Processing (NLP), focuses on languages which are spoken by majority populations and heavily standardised within one or more countries (Bender, 2009; Sawicki et al., 2023; Joshi et al., 2020). The field of NLP is thus reproducing the minoritisation of language varieties, and in the process reinforcing barriers of exclusion for diverse language communities. These include most local languages in highly linguistically diverse countries such as Indonesia (Aji et al., 2022), as well as various creoles which combine elements of settler-colonial languages elsewhere in the world (Lent et al., 2021). Additionally affected are varieties of national languages spoken by minoritised groups (Kantharuban et al., 2023), such as Māori English in New Zealand, Multicultural London English in the UK, and African-American English in the US. Speakers of minoritised varieties must typically switch to a majority variety when using language technologies, and when they do they can experience negative psychological impacts, such as feeling that technology wasn’t designed for them (Cunningham et al., 2024; Mengesha et al., 2021; Harrington et al., 2022; Wenzel et al., 2023).
@@
Machine learning-driven NLP is not typically oriented towards participatory processes, driven instead by the valorisation of efficiency (Birhane et al., 2022b) and fundamentally constrained by tensions between participation and scale (Young et al., 2024). NLP evaluations emphasise quantitative and decontextualised metrics (Hutchinson et al., 2022), while language dataset projects often reduce participation to crowdsourced platform work. These projects fail to create relationships which emphasise reciprocity and social justice (Sloane et al., 2022; Cooper and Zafiroglu, 2024), and risk perpetuating colonial histories, especially in the Global South, of exploitation by academic institutions, states and corporations (Markl, 2022). However participatory AI projects, on the other hand, are often challenged by a lack of shared understanding of good practices (Groves et al., 2023), and may risk “participation-washing” (Sloane et al., 2022). More generally, projects which aspire to participation often fail to afford decision-making power to participating communities (Arnstein, 1969), and risk co-optation when delivered through corporate structures (Asaro, 2000; Birhane et al., 2022a).
@@
In Australia, where this research takes place, of the 400+ ancestral Indigenous languages that existed before European colonisation (Bowern, 2023, p. 56), most are either extinct or critically endangered (Australian Government et al., 2020). This is due to linguicidal practices of colonialism, which at various times included governments forcibly separating Aboriginal children from their families and coercing Aboriginal pupils to speak English while forbidding them from speaking their languages (Rademaker, 2018). As a result, most Indigenous people in Australia speak Australian Aboriginal English, making it a strong encoder of ethno-cultural identity (Rodríguez Louro and Collard, 2024). [^1] Other contact languages in Australia include mixed languages such as Gurindji Kriol, Light Warlpiri and New Tiwi (Meakins, 2013). This indigenised variety of English integrates phonological, lexico-grammatical and discourse-pragmatic features from English and Aboriginal ancestral languages (see §2.1).
@@
To address these questions, this paper contributes the first detailed consideration of language technologies to support Australian Aboriginal English speakers. To our knowledge, this is the first paper to focus on technologies for this language community. We do so through a case study of an ongoing project intended to improve Automatic Speech Recognition (ASR) for Indigenous communities in Australia. In doing so, this paper follows previous NLP literature in considering a specific Indigenous language community and its context (e.g., Zhang et al., 2022; Teodorescu et al., 2022), with a focus on design considerations in order to avoid assuming that mainstream methods and goals are locally applicable (Noe and Kirshenbaum, 2024; Pine et al., 2024). As such, this paper contributes to our understanding of ethical and responsible development of technologies for minoritised communities, as well as to our understanding of how to design and build culturally sensitive technologies while incorporating participatory methods.
The paper begins with an introduction to Australian Aboriginal English, including a description of relevant linguistic and sociolinguistic factors (§2). We also present a brief survey of related work on Indigenous language technologies. Next, we discuss the opportunities and risks inherent in deploying ASR systems for this community, as well as opportunities and risks in projects building such technologies (§3). Building on this analysis, we explore strategies to mitigate risk, and facilitate opportunities, through a detailed discussion of the design considerations of a real-world project to design culturally appropriate ASR while integrating participatory methods (§4). In doing so, we provide a detailed discussion of some of the considerations that are critical for a project of this sort, including system requirements, data elicitation and transcription, system evaluation, and processes and frameworks for participation, and employment. Finally, we discuss further technological opportunities for supporting speakers of Australian Aboriginal English (§5).
@@ 2. Background and Related Work
When considering how technologies can support speakers of a specific Indigenous language variety, it is critical to have some basic knowledge of the speech community, the language variety, and their contexts (Zhang et al., 2022). In this section we introduce Australian Aboriginal English and its context, followed by a brief summary of prior work relevant to language technologies for Indigenous communities. This includes consideration of both ancestral and contact language varieties, and also of participatory methods in Indigenous language projects.
@@ 2.1. Australian Aboriginal English
Australian Aboriginal English is a contact variety of English used by Indigenous people in Australia (Rodríguez Louro and Collard, 2021a). It is particularly prevalent across southern Australia where ancestral languages are less widely spoken. In these contexts, Australian Aboriginal English is used to strengthen Indigenous ways of being and communicating (Eades, 2012a). Australian Aboriginal English has its origins in the product of forced contact between Indigenous people and the English-speaking settler-colonialists. The English spoken by the British was a powerful instrument used to strip Indigenous communities of their ancestral rights (Malcolm, 2000, p. 125). Australian Aboriginal English’s precursor was a mixed jargon viewed as ‘broken English’ by the British. By the time this jargon had transitioned to a novel pidgin (known as ‘New South Wales pidgin’ (Troy, 1994)), it began to travel outside present-day Sydney and across New South Wales. With the advent of the pastoral industry, the pidgin also spread north into Queensland and the Northern Territory, giving rise to Queensland Pidgin English and the Northern Australian creoles, including Roper River Kriol (Meakins, 2014, p. 371).
@@
Australian Aboriginal English has linguistic characteristics, including pronunciation, lexis, grammar, and discourse-pragmatics, that distinguish it from mainstream English (Eades, 2012a). Pronunciation features include: frequent /h/-dropping and /h/-addition; metathesis of ‘k’ and ‘s’ in the word ‘ask’, which is pronounced ‘aks’; and default reduction of the ‘ing’ verbal suffix to ‘in’. Grammatical features include double negatives as in ‘never done nothing’; and double subjects as in ‘My niece, she did it.’ Lexical features include lexical transfer such as blackfella, meaning ‘black person/fellow’, and borrowings from ancestral languages, such as dardy ‘good/good looking’ from Nyungar.
@@
Experts remain ambivalent as to whether Australian Aboriginal English represents a single ethnolect, or whether different varieties of Aboriginal English, or ‘Aboriginal Englishes’, exist across Australia (Dickson, 2019). Australian Aboriginal English is frequently placed on a continuum that ranges from ‘lighter’ acrolectal varieties (closer to standardised Australian English) to ‘heavier’ basilectal varieties which are closer to Kriol, spoken across northern Australia (Schultze-Berndt et al., 2013). The basilectal varieties are mostly unintelligible to standardised English speakers, but the acrolectal varieties are superficially similar to Australian English.
@@
Because Australian Aboriginal English is English based, it is sometimes socially stigmatised, erroneously classified by laypeople as an incorrect version of English. As a result, L1 users of Australian Aboriginal English are often discriminated against in schooling and legal settings (Harkins, 1994; Eades, 2012b), perpetuating invisibility of the variety and leading to the false perception that its speakers are uneducated and rude (Eades, 2013). Stigmatisation of Australian Aboriginal English is evident in the comment below, which was offered in response to an online article on the role of Australian Aboriginal English in Indigenous health communication (Rodríguez Louro and Collard, 2021).
@@
The commenter’s deliberate use of ‘Pidgin English’ reflects a deficit perspective that undermines the value of Australian Aboriginal English. Rather than embracing Australian Aboriginal English as a linguistic variety many Indigenous people use daily in Australia, this perspective portrays racialised peoples’ language varieties as ‘lacking’ vis-à-vis mainstream English (Hudley et al., 2020, p. e214).
These negative attitudes are exacerbated because Australian Aboriginal English is primarily oral, with no nationally adopted guidelines as to how it should be written. Additionally, the orthographic representation of Australian Aboriginal English varies across locales, reflecting the local linguistic biographies and practices of distinct Indigenous groups. For example, the name of the Traditional Owners of the region comprising the southwest of Western Australia, as well as their language, is variably spelled ‘Nyungar’, ‘Noongar’, or ‘Nyoongar’ depending on community members’ preferences. [^2] Collard is Nyungar so we prefer this spelling. Rather than a weakness, the lack of standardising orthographies for Australian Aboriginal English reflects Indigenous people’s prerogative to differentiate—rather than homogenise—how Australian Aboriginal English is used across Indigenous Country.
Indigenous people tend to find the use of direct questions awkward or offensive (Rodríguez Louro and Collard, 2024), [^3] With some exceptions, including teaching contexts where instructors asking questions is heavily routinised (Moses et al., 2008). and these pragmatic considerations continue in Australian Aboriginal English. Direct questions are also likely to elicit misleading information; Indigenous people have been known to answer ‘yes’ or agree with non-Indigenous people even if the question is not fully understood—a phenomenon known as ‘gratuitous concurrence’ (Liberman, 1985, p. 137). To avoid these limitations, researchers have used ‘yarning’, an Indigenous cultural form of storytelling and conversation, to lead unstructured data collection with Indigenous communities who speak Australian Aboriginal English (Bessarab and Ng’andu, 2010; Rodríguez Louro and Collard, 2021b).
@@ 2.3. Technologies for contact language varieties
The NLP literature has to-date had marginal concern with supporting Indigenous contact language communities, such as Hawaiian Pidgin, Australian Kriol, Tok Pisin, Chavacano, etc. For example, Nigerian Pidgin has over 100 million speakers (Lin et al., 2024), yet is poorly supported by technology. This seems to be due to a combination of reasons, including the stigmatisation of contact languages (Siegel, 1999; Lent et al., 2024) and a belief that ancestral languages are more important in terms of global heritage. However Indigenous peoples are not museum artifacts or relics of an ‘ancient’ era (Gaby and Woods, 2020; Hill, 2002); their cultures and societies are dynamic and evolving.
@@ 2.4. Culture, and participatory methods, in language research
Recently, the NLP community has demonstrated increasing interest in the topic of culture, with, e.g., three recent workshops on Cross-cultural Considerations in NLP (Dev et al., 2023; Prabhakaran et al., 2024, 2025). This interest has been spurred by the capabilities of Large Language Models (LLMs) with regards to linguistic form, while also demonstrating lack of cultural nuance (Hershcovich et al., 2022; Naous et al., 2024; Li et al., 2024). Threads of research have focused on measuring cultural biases (Naous et al., 2024; Tao et al., 2024) and expression of cultural values by LLMs (Arora et al., 2023; Durmus et al., 2023; Ramezani and Xu, 2023).
@@ 2.5. Projects with Indigenous communities
Indigenous language varieties are typically differentiated functionally within a multilingual community, and technologies may be more useful for some than others (Doğruöz and Sitaram, 2022; Bird, 2022; Lent et al., 2022). Technologists should understand the technology needs of Indigenous communities, which may be distinct from those of non-Indigenous users (Liu et al., 2022; Cooper et al., 2024).